OpenStudy (owlcoffee):

Find the equations of the lines tangent to x^2+y^2-x+3y-2=0 with the direction of the line 2x+2y-5=0.

OpenStudy (amistre64):

define y' implicitly, and equate it to the slope of the given line

OpenStudy (owlcoffee):

You mean to first put it in ordinary form to deduce it's slope?

OpenStudy (amistre64):

the line part, sure. that is one way to find its slope.

OpenStudy (amistre64):

an implicit rundown for: x^2+y^2-x+3y-2=0 2x + 2y y' -1+ 3y' = 0 solve for y' and equate it with the slope of the line

OpenStudy (owlcoffee):

by Equating, do you mean evaluating?

OpenStudy (amistre64):

hmm, maybe. depends on where youre from i spose. the slope of the given line is -1, right? and y' defines the slope of the line at any given point on the curve. when y' = -1, we know the point at which the line is tangent to the curve; and can use the pont slope form of a line to define the tangent line equations with

OpenStudy (amistre64):

2x + 2y y' -1+ 3y' = 0 y' (2y + 3) = 1-2x y' = (1-2x)/(2y+3) (1-2x)/(2y+3) = -1 1 - 2x = -2y - 3

OpenStudy (amistre64):

using y = x-2 as a solution; sub it into the curve x^2+y^2-x+3y-2=0 x^2+(x-2)^2-x+3(x-2)-2=0; and solve for x (and y as a result).

OpenStudy (amistre64):

i have class starting so I have to leave ... good luck

OpenStudy (owlcoffee):

Thank you for your time, I am very grateful. Good luck with your class ~

OpenStudy (amistre64):

in solving for x x^2+(x-2)^2-x+3(x-2)-2=0 x^2+(x^2-4x+4)-x+(3x-6)-2=0 2x^2 - 2x - 4=0 2 (x^2 - x - 2)=0 we see that x = -1, or 2 giving us: y = -1-2, or 2-2 to construct the tangent lines with a slope of -1

OpenStudy (owlcoffee):

I see. I'll try something, hhow about using the line family that go through that point, and consider the lines tangent to the circle. All will be paralell so ther shouldn't be any trouble.

OpenStudy (amistre64):

since: x^2+y^2-x+3y-2=0 , is a circle, you could perp the slope of the given line to 1, and attach it to the center of the circle and determine where it crosses the circle at as well. But that is more of a geometric interpretation than a calculus based approach :)

OpenStudy (owlcoffee):

basically the tangent lines and the line that crosses the center should have the same slope. using the line family would give the same slope as if using implicit differentiation. Again, all those are methods hehe...

OpenStudy (amistre64):

indeed :)